# Domain Generalization in LiDAR Semantic Segmentation Leveraged by Density Discriminative Feature Embedding

### v1

@@ IV-D Ablation Studies
To validate the contributions of each component in our method, we conduct ablation studies focusing on point-voxel encoding, the density-aware embedding module, density soft clipping, and density augmentation. These experiments align with the LiDomAug [12] settings.
IV-D1 Density discriminative feature embedding
In Tab. VII, which extends Tab. IV, we present an ablation study using the experimental setup proposed by LiDomAug [12] and based on MinkowskiNet42. This study highlights the key components of our DDFE module: (a) point-voxel encoding, (b) density-aware embedding module, and (c) density clipping. Additionally, we evaluate (d) the effect of our density augmentation method. When all these components are utilized, we observe a notable performance increase of +13.5% in the (KN) scenario and +33.1% in the (NK) scenario, compared to the baseline. The density-aware embedding module emerges as a particularly impactful factor, closely followed by the point-voxel feature encoding, which underscores its significance. The density soft clipping method proves its worth by enhancing the model’s adaptability to new density patterns, while the improvements due to the density clipping technique underscore the challenges models face with unfamiliar density distributions.
IV-D2 Data augmentation
In Tab. VIII, we compare our density augmentation method with existing data augmentation techniques. To adapt the Mix3D method to our DDFE module, we modify the density value mixing process as outlined in Alg. 2, specifically at line #10. Our analysis reveals that aligning the training dataset with the target density characteristics is crucial for performance improvement. This is evident with Beam drop augmentation in the density decreasing (KN) scenarios and Mix3D in density increasing (NK) scenarios, where each method leads to substantial gains. Conversely, scenarios where this alignment is not achieved see limited performance improvements. Furthermore, our Enhanced-Mix3D (E-Mix3D) outperforms the original Mix3D during the transition from nuScenes to KITTI, benefiting from a more varied density distribution. Our approach, which combines E-Mix3D with beam drop augmentation, not only surpasses methods without augmentation, showing significant performance improvements of +8.4% and +10.8%, but also yields additional gains of +1.8% and +4.0% over the standard Mix3D and beam drop combination. This highlights the effectiveness of our density augmentation strategy in enhancing performance across different scenarios.
| Method | KN | NK |
| w/o aug | 46.2 | 41.8 |
| Beam drop [13] | 49.1 (+6.3%) | 42.7 (+2.2%) |
| Mix3D [37] | 46.7 (+1.1%) | 44.0 (+5.3%) |
| Mix3D + Beam drop | 49.2 (+6.5%) | 44.5 (+6.5%) |
| E-Mix3D | 48.7 (+5.4%) | 45.4 (+8.6%) |
| E-Mix3D + Beam drop | 50.1 (+8.4%) | 46.3 (+10.8%) |
IV-D3 Feature similarity
To analyze the effects of DDFE, we investigate the feature similarity between the source and unseen dataset as shown in Fig. 7. For our analysis, we consider the input voxel features of the 3D backbone model generated by the DDFE module. We begin our investigation with a baseline model that employs a point-voxel feature encoding method. This is then compared against our method integrated a density-aware embedding module and a density soft clipping. Both of these models are trained on the nuScenes dataset and then tested on the Waymo dataset, which served as the unseen dataset. We compute similarity values based on the L2 distance between the average features over specified distance intervals (e.g., 0-5m, …, 45-50m) of each dataset. From the visualization in Fig. 7, the baseline model consistently demonstrates higher feature similarity at similar distances, regardless of density. This can be attributed to the feature encoding process that directly embeds the 3D coordinates of points. On the other hand, the model that combines the baseline with the density-aware embedding module with soft clipping shows pronounced feature similarity within consistent density distributions, even when making comparisons between the source and unseen datasets.
V Conclusion
@@
In this paper, we introduce a new perspective for domain generalization, utilizing the inherent property of LiDAR where density varies with distance. Based on this perspective, we propose a density discriminative feature embedding module called DDFE, which enhances the ability to discern densities by estimating expected densities from LiDAR beams. As a result, it improves the generalization performance, especially when transitioning between the source dataset and previously unseen datasets. In addition to the core methodology, we also present a simple and effective data augmentation technique. This method amplifies the density spectrum of a point cloud without relying on ego-motion or sequential labeled data. Extensive experiments on the SemanticKITTI, Waymo, and nuScenes datasets demonstrate the effectiveness of the proposed method in elevating the domain generalization performance of LiDAR semantic segmentation. In future work, we aim to extend domain generalization for other LiDAR-based perception tasks such as 3D object detection by leveraging object-wise density discriminative features in conjunction with point-wise density embedding.

### v2

@@ 4.4 Ablation Studies
To validate the impact of individual components within our method, we conduct ablation studies. These studies concentrate on the evaluation of point-voxel encoding, the density-aware embedding module, density soft clipping, and density augmentation, providing insights into the significance of each element in enhancing the overall methodology.
4.4.1 Analysis of the individual components of our method
In our comprehensive ablation study in Tab. 4, we thoroughly investigate the individual components of our method. We use the experimental setup proposed by LiDomAug [29] and employ MinkowskiNet [7] as the backbone. This study dissects the impact of our DDFE, which includes (a) point-voxel encoding, (b) density-aware embedding module, and (c) density clipping. Additionally, it examines (d) the impact of our density augmentation method.
The integration of all these elements results in a significant performance uplift +13.5% in the (KN) scenario and +33.1% in the (NK) scenario over the baseline model. Notably, the density-aware embedding module stands out for its substantial effect, closely followed by the point-voxel encoding, highlighting its critical role. Furthermore, the study validates the utility of density soft clipping in enhancing model flexibility towards novel density configurations. Meanwhile, the gains observed with density clipping highlight the inherent challenge in adapting to unfamiliar density landscapes, emphasizing the nuanced contributions of each component towards achieving robust domain generalization. Lastly, incorporating density augmentation into DDFE leads to further enhancements, with an additional 8.4% performance increase in the (KN) scenario and 10.8% in the (NK) scenario. This demonstrates the effectiveness of the proposed density augmentation scheme in significantly boosting the performance of DDFE.
4.4.2 Augmentation
Tab. 5 shows the performance comparison when various augmentation methods are applied to the proposed DDFE. The experimental setup follows DGLSS [16], and the baseline involves applying the proposed DDFE to MinkowskiNet [7]. The proposed enhanced-Mix3D shows a 2.3% performance improvement over the original Mix3D [22] when trained on SemanticKITTI and tested on nuScenes (KN), and a 3.6% improvement when trained on nuScenes and tested on SemanticKITTI (NK). In conjunction with the beam sampling method, it achieves a 7.6% performance improvement in (KN) and 11.2% in (NK) compared to the baseline.
| Method | D. V. | KN | NK |
| DDFE | 45.9 | 41.8 | |
| + PolarMix [40] | 48.2 (+5.0%) | 42.3 (+1.2%) | |
| + B. S. [38] | ✓ | 49.1 (+7.0%) | 42.2 (+0.9%) |
| + Mix3D [22] | ✓ | 47.4 (+3.2%) | 44.8 (+7.1%) |
| + Mix3D + B. S. [38] | ✓ | 49.1 (+7.0%) | 44.2 (+5.7%) |
| + E-Mix3D | ✓ | 48.5 (+5.6%) | 46.4 (+11.0%) |
| + E-Mix3D + B. S. [38] | ✓ | 49.4 (+7.6%) | 46.5 (+11.2%) |
4.4.3 Analysis of Density Discriminative Feature Embedding (DDFE)
We delve into the effectiveness of the DDFE module by examining the feature similarity between the source dataset, nuScenes, and the unseen dataset, Waymo, as depicted in Fig. 4. Our focus is on the input voxel features of the 3D backbone model generated by the DDFE. We initiate our analysis by setting a baseline that utilizes a point-voxel feature encoding method for comparison. Following this, we proceed to assess our enhanced method, which integrates a density-aware embedding module along with density soft clipping, to understand its impact on bridging the domain gap. Similarity metrics are derived from the L2 distance between averaged features across specified distance intervals (e.g., 0-5m, …, 45-50m) for each dataset. According to Fig. 4, the baseline consistently shows higher feature similarity across distances, influenced by its direct embedding of 3D point coordinates. Conversely, our integrated model with density-aware embedding and soft clipping significantly aligns feature similarity across consistent density distributions. It demonstrates the efficacy of the proposed modules in adapting to differences between the source and unseen datasets.
4.4.4 Computational Cost
The inference time of our method for processing a single frame with the nuScenes is 44ms on an NVIDIA RTX 3090. The inclusion of our DDFE module into the MinkNet42 architecture accounts for an extra 8ms of this computation time, while the base MinkNet42 architecture alone requires 36ms. This integration modestly elevates the model’s complexity, introducing approximately 23.8k parameters (+0.06%). Furthermore, the application of density augmentation during training adds an extra 60ms per scan on the nuScenes.
5 Conclusion
@@
In this paper, we introduce a new perspective for domain generalization of LiDAR semantic segmentation by exploiting the concept of density diversity within a source domain. Based on this perspective, we propose the Density Discriminative Feature Embedding (DDFE) module. DDFE is designed to incorporate expected densities derived from LiDAR beams into a density-aware feature space, significantly enhancing the model’s capability to distinguish between different densities. This novel approach improves the adaptability and accuracy of LiDAR semantic segmentation, enabling them to perform effectively across diverse domains, including those previously unseen. In addition to the core methodology, we also present a simple and effective data augmentation technique that extends the density spectrum of the source data. Extensive experiments on the SemanticKITTI, Waymo, and nuScenes datasets demonstrate the effectiveness of the proposed method in elevating the domain generalization performance of LiDAR semantic segmentation. A limitation of this study is its exclusive focus on semantic segmentation tasks. Thus, in future work, we plan to extend our domain generalization method to encompass other LiDAR-based perception tasks, such as 3D object detection. This extension can involve utilizing object-centric density discriminative features alongside point-wise density embedding.
Acknowledgements
This work was supported by the Digital Innovation Hub project supervised by the Daegu Digital Innovation Promotion Agency(DIP) grant funded by the Korea government(MSIT and Daegu Metropolitan City) in 2023(DBSD1-02, Vision picking system for the logistics industry based on artificial intelligence object recognition) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00210908).